OpenStudy (anonymous):

Can someone help me on my math work? I'm not sure how to do it. You don't have to help me on every question, but if you could do a couple with that would be great, just so i know how to do the rest. I have a midterm coming up and I want to make sure i'm doing it correctly.

OpenStudy (readergirl12):

I'll try...What grade are you in?

OpenStudy (anonymous):

I'm in 10th grade.

OpenStudy (readergirl12):

Hm...Okay give me a moment.

OpenStudy (anonymous):

okay

OpenStudy (zpupster):

hey were you helped

OpenStudy (zpupster):

sorry now i have to run.

OpenStudy (kobeni-chan):

Do you still need help? If not, please close the question :)

OpenStudy (anonymous):

Yes. I still need some help.

OpenStudy (whpalmer4):

which part do you need help with?

OpenStudy (whpalmer4):

Looking at your first page, a common task seems to be to find out if the lines are parallel or perpendicular. Two lines which are parallel will have an identical slope. Two lines which are perpendicular will have slopes such that if you multiply them together, the result is \(-1\). The only case where this does not hold true is if your lines are parallel to the x and y axes, in which case their equations will be of the form \[x=c\]and\[y=d\]where \(c,d\) are constants. A vertical line has an undefined slope, and a horizontal line has a slope of \(0\), so you cannot multiply them together to get \(-1\). If you rearrange your equation for the line to be in what is called slope-intercept form: \[y = mx + b\]\(m\) is the slope and \(b\) is the \(y\)-intercept value It is easy to compare two equations in this form and determine if they are parallel (the number in front of the \(x\) will be the same\) or perpendicular (multiply the number in front of the \(x\) from each equation, if the product is \(-1\), they are perpendicular).

OpenStudy (anonymous):

Can you show me how to do the first problem where you have to state the slope and tell whether its parallel perpendicular or neither? So I can I have an example to follow.

OpenStudy (whpalmer4):

I thought I already did :-) \[y = \frac{4}{3}x+5\]Compare that with the slope-intercept form:\[y=mx+b\]what is the value of \(m\) and what is the value of \(b\)? The value of \(m\) is the slope.

OpenStudy (whpalmer4):

the other line has equation \[y = -\frac{3}{4}x-1\]Again, compare with \[y=mx+b\]What is the value of \(m\)? Is it equal to the \(m\) in the previous equation? If it is, they are parallel. If it is not, multiply them together. Do you get \(-1\)? If so, they are perpendicular. That's all there is to it.

OpenStudy (anonymous):

Oh, okay. Thank you! I have a midterm next week, and I'm so nervous.